# Conformal Tradeoffs: Guarantees Beyond Coverage

### v1

@@ Abstract
Deployed conformal predictors are long-lived decision infrastructure operating over finite operational windows. The real-world question is not merely “Does the true label lie in the prediction set at the target rate?”, i.e. marginal coverage, but “How often does the deployed system commit versus defer? What error exposure does it induce when it acts? How do these operational rates trade off?” Marginal coverage does not summarize these deployment-facing quantities: the same calibrated thresholds can yield very different operational profiles depending on score geometry.
This paper provides a framework for operational certification and planning beyond coverage. We make three contributions. (1) Small-Sample Beta Correction (SSBC): We invert the exact finite-sample Beta/rank law for split conformal to map a user coverage request $(\alpha^{\star},\delta)$ into a concrete calibrated grid point with PAC-style semantics, providing explicit finite-window coverage guarantees for the deployed rule. (2) Calibrate-and-Audit: Because no distribution-free pivot exists for rates beyond coverage, we introduce a two-stage design: an independent audit set yields a reusable region–label table that supports certified finite-window predictive envelopes (Binomial/Beta–Binomial) for operational quantities—commitment frequency, deferral, decisive error exposure, commit purity—or related quantities via linear projection, without committing to a specified scalar objective. (3) Geometric characterization: We expose the feasibility constraints, regime boundaries (hedging vs. rejection), and cost-coherence conditions induced by a fixed conformal partition, explaining why operational rates are coupled and how calibration navigation trades them off.
The output is an auditable operational menu: given a scoring model, the framework traces the Pareto frontier of attainable operational profiles across calibration settings and attaches finite-window uncertainty envelopes to each regime. We demonstrate the approach on Tox21 toxicity prediction (12 endpoints) and aqueous solubility screening of lipophilic drug candidates via the AquaSolDB dataset.
@@ 1 Introduction: deployment-facing conformal prediction
Many classifiers are deployed as long-lived decision infrastructure rather than one-off prediction engines. A single trained scoring model is integrated into downstream workflows and reused over finite operational windows. In this regime, the deployed object is not the score function alone, but the rule: (scoring model, one-time calibration procedure, and downstream action convention). The observables that matter to stakeholders are finite-window operational rates: how often the system commits versus defers/abstains, and—conditional on commitment—the decisive error exposure and purity induced by attached actions (cf. rejection options and selective classification (Chow, 1970)). These realized rates drive resource planning, safety and compliance, and near-term impact.
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) is a natural starting point for deployment-facing calibration: it sets score thresholds on held-out data and yields finite-sample, distribution-free coverage under exchangeability (Papadopoulos et al., 2002). Coverage, however, is not an operational specification. Once a deployment convention is fixed (e.g., “commit on singletons, defer on hedges or empty sets”), the realized operational profile depends on how calibration choices partition score space and how mass and labels populate the induced regions. As a result, two conformal rules with identical marginal coverage can behave very differently in deployment.
@@ 1.1 The gap: coverage does not determine operational behavior
Standard conformal guarantees that the true label lies in the prediction set at the target coverage rate (under exchangeability). But coverage alone does not determine the operational profile of a deployed conformal rule. In particular, different calibrated predictors with the same nominal coverage can exhibit markedly different:
@@ 1.2 Our approach: calibrate, expose geometry, and audit once
We study a single deployed split-conformal rule through a calibration-conditional, certification-style lens. A calibration choice fixes cutpoints on the score axis, inducing a finite partition of score/output space into regions with set-valued outcomes (e.g., singleton, hedge, abstain). The key device in this paper is to make this intermediate geometry explicit and auditable: once the region partition is fixed, many policy- and Key Performance Indicator (KPI)-specific rates are obtained by simple projection from the joint region–label masses. This supports trade-off analysis across calibration settings and enables distribution-free uncertainty statements about future realized finite-window rates for the deployed rule.
Figure 1 previews the main object produced by this paper: an auditable operational menu mapping calibration settings to certified finite-window operational profiles, together with a Pareto filter that highlights nondominated operating regimes.
Figure 1: Operational menu induced by a calibrated conformal rule on a synthetic model. Left: a calibration choice fixes score cutpoints that partition score space into set-valued outcomes (e.g., singleton, hedge, abstain) and associated region–label probabilities. A deployment convention determines which outcomes trigger commitment versus deferral. Right: sweeping calibration settings traces the attainable set of selected operational rates; the oriented Pareto frontier highlights nondominated operating regimes.
@@ 1.3 Contributions
This paper makes operational quantities first-class objects of certification and exploration for deployed conformal predictors. Our contributions are:
@@
We invert the exact finite-sample rank/Beta law for split conformal (Marques, 2025) to map a user request $(\alpha^{\star},\delta)$ (e.g., “90% coverage with 90% confidence”) to the least conservative split-conformal grid point satisfying a PAC-style tail constraint:
@@
Operational rates such as commitment, deferral, and decisive error exposure are not rank-determined and therefore do not admit a conformal “coverage-style” pivot. We introduce a two-stage design:
@@
  • Audit: use an independent exchangeable set $\mathcal{D}_{\mathrm{audit}}$ to estimate the joint region–label table $\{p_{r,y}\}$.
The joint table is a reusable sufficient statistic: a wide class of operational KPIs are linear projections (or monotone transforms) of $\{p_{r,y}\}$. This model, via an induced Binomial/Beta–Binomial sampling model of events of interest, yields exact finite-window predictive envelopes for future realized operational rates, enabling certified policy comparison across many calibration settings without re-auditing. We also provide a conservative leave-one-out proxy for single-sample settings (Appendix D).
@@
Together, these results explain why some desired operational profiles are infeasible, and provide better guidance for practitioners to navigate emergent trade-offs induced by conformal calibration choices.
@@ 1.4 Positioning and related work
Conformal prediction constructs set-valued predictors with finite-sample coverage guarantees under exchangeability (Vovk et al., 2005; Shafer and Vovk, 2008; Angelopoulos and Bates, 2023). Conformal risk control (CRC) generalizes this viewpoint by selecting thresholds to satisfy user-specified scalar risk constraints under related assumptions (Angelopoulos et al., 2024; Bates et al., 2021), and extensions address covariate shift via reweighting or adaptation (Tibshirani et al., 2019; Fannjiang et al., 2022). Recent work also connects conformal methods to downstream decision-making: conformal decision theory calibrates decision rules (often online) to control realized risk (Lekeufack et al., 2023), and decision-theoretic foundations motivate conformal sets as uncertainty objects for risk-averse agents (Kiyani et al., 2025).
Our focus is complementary. Rather than guaranteeing coverage or a single scalar risk functional, we certify finite-window operational rate vectors for a fixed deployed rule and support multi-objective exploration without committing to a scalar objective. The audited primitive is the intermediate geometry induced by calibration—the region partition and its joint region–label table—from which many deployment policies and KPIs follow by projection.
Closest in spirit to our “menu” viewpoint is the inverse CRC of Zhou and Zhu (2025), who trace certified miscoverage–regret trade-offs across robustness levels for predict-then-optimize pipelines; they audit coverage and regret as the downstream objects of interest. In our work, however, we target an earlier deployment stage, where the cost model and operational targets are not yet fixed. At that stage we audit and certify operational quantities—commit and deferral rates, error exposure, purity, and workload—and their 95% predictive envelopes that directly govern resource usage and risk over finite windows.
@@
A different line of work optimizes efficiency by selecting, from a family of candidate models, the conformal predictor that minimizes a set-size functional (e.g., interval length/volume, or prediction-set cardinality in classification) subject to validity constraints (Yang and Kuchibhotla, 2021). This is well-motivated when downstream utility is monotone in set size. Our objective differs: in many deployments, smaller cardinality does not necessarily imply higher utility, nor lower uncertainty in the induced region-conditional label frequencies. In particular, aggressive singleton behavior can reduce average set size while increasing “decisive error exposure” and degrading the operational profile that stakeholders ultimately care about. When organizational priorities are still being determined, a set size proxy might not be informative enough to guide decision-making. Accordingly, we treat conformal sets as instruments whose value is expressed through a vector of auditable KPIs, and we certify the feasible trade-offs among these rates in finite windows.
@@ 1.5 Roadmap
Section 2 formalizes the calibration-conditional viewpoint for a fixed deployed object and defines the region–label table as the auditable primitive. Section 3 presents SSBC and the Calibrate-and-Audit framework, yielding certified finite-window envelopes for operational rates and Pareto-based trade-off navigation. Section 4 characterizes regime boundaries and feasibility constraints induced by the conformal partition. Section 5 validates the framework empirically on Tox21 toxicity prediction and aqueous solubility prediction, and Section 6 discusses limitations and extensions (multiclass, online adaptation, and distribution shift).
@@ 2 Setting and notation: calibration-conditional viewpoint
We study a single deployed predictor produced by (i) training a scoring model on a training set and then treating it as fixed, followed by (ii) a one-time split conformal calibration on an exchangeable calibration sample. Accordingly, all randomness in our analysis is over the one-time calibration draw and future deployment data; we do not analyze variability due to retraining or repeated recalibration. Later, to make operational rates beyond coverage auditable with certified finite-sample guarantees, we introduce a separate exchangeable audit set held out alongside the calibration sample.
@@ 2.1 Data splits and exchangeability assumptions
Let $\mathcal{D}_{\mathrm{train}}$ denote a training dataset used to fit a base scoring model, for instance a probabilistic classifier (Goodfellow et al., 2016). After training, the resulting score function is treated as fixed.
@@
The calibration sample $\mathcal{D}_{\mathrm{cal}}$ is used once to set the deployed rule’s thresholds; the audit sample $\mathcal{D}_{\mathrm{audit}}$ is held out (e.g., by splitting the calibration pool) and reserved for measuring operational quantities after thresholds are fixed, and is not used to choose thresholds. This approach is equivalent to the double calibration set approach of Yang and Kuchibhotla (2021).
@@ 2.2 Scores and calibration thresholds
Let $\mathcal{Y}=\{1,\ldots,K\}$, and let $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ denote a nonconformity (score) function (Shafer and Vovk, 2008; Lei et al., 2018). A common choice for probabilistic-like classifiers is $s(x,y)=1-P(y\mid x)$. While this specific form is not required for the present section, it is used in the examples throughout the paper.
@@
The key operational point is that, given $(\mathcal{D}_{\mathrm{cal}},\alpha^{\star})$, calibration produces a vector of class-conditional thresholds $\tau=(\tau_{1},\ldots,\tau_{K})$, and these thresholds are then held fixed throughout deployment.
@@ 2.3 Regions, observability, and policies
and (optionally class-conditional) thresholds
@@
Thus calibration fixes a region map $R_{\tau}:\mathcal{X}\to\{0,1\}^{K}$, i.e., a finite partition of score space used throughout deployment. The map $R_{\tau}$ is determined entirely by $\tau$. However, the induced distribution of class labels $Y$ given the region $R_{\tau}(X)$—which regions carry mass and how much—depends on the deployment data-generating process. This discrete region partitioning is our decision and exploration interface.
Figure 2: Region definitions and region mass under different data regimes. (A) Thresholds $\tau_{0},\tau_{1}$ induce a finite region partition $R_{\tau}(x)$ independent of data and policy. (B–C) Under probability-normalized scores, data support is restricted to the diagonal, determining which regions carry nonzero mass for different threshold configurations. (D) With unconstrained scores, all regions may carry mass. No deployment policy is applied in this figure.
@@
The standard split conformal predictor corresponds to the set inclusion policy (Romano et al., 2019; Barber et al., 2021)
@@ 2.4 Auditing primitives: region–label tables and policy projections
Conditional on $\mathcal{D}_{\mathrm{cal}}$, the deployed rule is fixed, and the primitive observable for auditing is the pair $(R_{\tau(\theta)}(X),Y)$ on an exchangeable labeled split. Two representations are useful: the full region–label table, and KPI indicators obtained as its projections. Here $\theta\in\Theta$ denotes a calibration setting that determines thresholds $\tau(\theta)$ and hence the region map $R_{\tau(\theta)}$.
(A) Region–label joint table.
@@
Note that the region map $R_{\tau(\theta)}$ is fixed from $\mathcal{D}_{\mathrm{cal}}$, while the counts $K^{\mathrm{audit}}_{r,y}(\theta)$ are computed on $\mathcal{D}_{\mathrm{audit}}$ only.
Because $R_{\tau(\theta)}$ is a finite partition of score space, the region–label events $(R_{\tau(\theta)}(X)=r,\ Y=y)$ form a finite partition of the sample space. Consequently, for every $\theta$,
@@
Moreover, the label marginals are fixed by the deployment data-generating process:
$\sum_{r\in\mathcal{R}}p_{r,y}(\theta)=\mathbb{P}(Y=y\mid\mathcal{D}_{\mathrm{cal}}),\qquad\sum_{r\in\mathcal{R}}\widehat{p}^{\,\mathrm{audit}}_{r,y}(\theta)=\widehat{\mathbb{P}}^{\,\mathrm{audit}}(Y=y),$
so varying $\theta$ reallocates mass across regions within each label. These conservation laws are the basic source of geometric coupling and explain why sweeping $\theta$ produces a constrained set of rate vectors rather than filling an ambient volume.
@@
Thus the region–label table is a reusable sufficient audit summary: audit it once, then compute policy-level KPIs by projection. The same structure extends to monotone transformations of linear KPIs (e.g., ratios or log-odds), since the region–label table remains the sufficient summary.
@@
$P(\theta)=\begin{array}[]{c|ccc}&y=0&y=1&\hat{C}_{\pi_{\mathrm{SI}}}(X)\\ \hline\cr r_{10}=(1,0)&\mathbf{p_{10,0}(\theta)}&p_{10,1}(\theta)&\{0\}\\ r_{11}=(1,1)&\mathbf{p_{11,0}(\theta)}&\mathbf{p_{11,1}(\theta)}&\{0,1\}\\ r_{01}=(0,1)&p_{01,0}(\theta)&\mathbf{p_{01,1}(\theta)}&\{1\}\\ r_{00}=(0,0)&p_{00,0}(\theta)&p_{00,1}(\theta)&\varnothing\end{array}\qquad\text{with}\qquad\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}p_{r,y}(\theta)=1.$
@@
The same “sum selected cells” structure applies to abstention/deferral, decisiveness, and decisive error exposure under any fixed policy. This is the sense in which operational quantities become first-class: it yields a single auditable summary from which many KPIs can be computed (and later renegotiated) by projection, while making the conservation constraints that drive trade-offs explicit.

### v2

@@ Abstract
Deployed conformal predictors are long-lived decision infrastructure reused over finite operational windows. In practice, stakeholders care not only about marginal coverage, but also about operational quantities: how often the system commits versus defers, and what error exposure it induces when it acts. These deployment-facing quantities are not determined by coverage alone: identical calibrated thresholds can yield markedly different operational profiles depending on score geometry. We develop tools for operational certification and planning beyond coverage for split conformal prediction. First, Small-Sample Beta Correction (SSBC) inverts the exact finite-sample rank/Beta law to map a user request $(\alpha^{\star},\delta)$ to a concrete calibration grid point with PAC-style semantics, yielding explicit finite-window coverage guarantees for a reused deployed rule. Second, because no distribution-free pivot exists beyond coverage, we propose Calibrate-and-Audit: an independent audit set supports certified finite-window predictive envelopes (Binomial/Beta-Binomial) for key operational quantities—commitment frequency, deferral, and decisive error exposure—and related metrics via linear projection, without committing to a scalar objective. Third, we give a geometric characterization of the feasibility constraints and regime boundaries induced by a fixed conformal partition, clarifying why operational quantities are coupled and how calibration navigation trades them off. The result is an operational menu rthat traces attainable operational profiles (Pareto trade-offs) and attach finite-window uncertainty envelopes to each regime. We illustrate the approach on benchmark molecular toxicity and aqueous solubility datasets.
@@ 1 Introduction: deployment-facing conformal prediction
Many classifiers are deployed as long-lived decision infrastructure rather than one-off prediction engines. A single trained scoring model is integrated into downstream workflows and reused over finite operational windows. In this regime, the deployed object is not the score function alone, but the combined rule tying together the scoring model, the one-time calibration procedure, and the downstream action convention. The observables that matter to stakeholders are finite-window operational rates: how often the system commits versus defers/abstains, and conditional on commitment the decisive error exposure and purity induced by attached actions (cf. rejection options and selective classification (Chow, 1970)). These realized rates drive resource planning, safety and compliance, and near-term impact.
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) is a natural starting point for deployment-facing calibration: it sets score thresholds on held-out data and yields finite-sample, distribution-free coverage under exchangeability (Papadopoulos et al., 2002). Coverage alone, however, is not an operational specification that captures the full range of deployment-facing quantities (Hullman et al., 2025). Once a deployment convention is fixed (e.g., “commit on singletons, defer on hedges or empty sets”), the realized operational profile depends on how calibration choices partition score space and how mass and labels populate the induced regions. As a result, two conformal rules with identical marginal coverage can behave very differently in deployment.
@@ 1.1 The gap: coverage does not determine operational behavior
Standard conformal guarantees that the true label lies in the prediction set at the target coverage rate (under exchangeability). But coverage alone does not determine the operational profile of a deployed conformal rule. In particular, different calibrated predictors with the same nominal coverage can exhibit markedly different operational characteristics. Examples include:
@@ 1.2 Our approach: calibrate, expose geometry, and audit once
We study a single deployed split-conformal rule through a calibration-conditional, certification-style lens. A calibration choice fixes cutpoints on the score axis, inducing a finite partition of score/output space into regions with set-valued outcomes (e.g., singleton, hedge, abstain). The key device in this paper is to make this intermediate geometry explicit and auditable: once the region partition is fixed, many policy- and Key Performance Indicator (KPI)-specific rates are obtained by simple projection from the induced joint region–class label distribution. Our development focuses on binary classification. Even in this simplest nontrivial setting, the induced geometry already reveals the main operational trade-offs. Within this setting, we can study trade-offs across calibration settings and derive distribution-free uncertainty statements about future realized finite-window rates for the deployed rule.
Figure 1 previews the main object produced by this paper: the map from calibration choices to the induced joint region–class label distribution. This representation can then be interrogated through Pareto analysis to expose operational trade-offs before a price point or consequence cost has been specified.
Figure 1: Operational menu induced by a calibrated conformal rule on a synthetic model. Left: a calibration choice fixes score cutpoints that partition score space into set-valued outcomes (e.g., singleton, hedge, abstain) and associated region–class label probabilities. A deployment convention determines which outcomes trigger commitment versus deferral. Right: sweeping calibration settings traces the attainable set of selected operational rates; the oriented Pareto frontier highlights nondominated operating regimes.
@@ 1.3 Contributions
This paper studies operational quantities as explicit objects of exploration and certification for deployed conformal predictors. Our contributions are:
@@
We invert the exact finite-sample rank/Beta law for split conformal (Marques, 2025; Angelopoulos and Bates, 2023) to map a user request $(\alpha^{\star},\delta)$ (e.g., “at least 90% coverage with at least 90% confidence”) to the least conservative split-conformal grid point satisfying a Probabily Approximately Correct (PAC) tail constraint:
@@
Operational rates such as commitment, deferral, and decisive error exposure are not rank-determined like coverage and therefore do not admit a conformal “coverage-style” pivot that determines their distribution. We therefore introduce a two-stage design for studying these quantities:
@@
  • Audit: use an independent exchangeable set $\mathcal{D}_{\mathrm{audit}}$ to estimate the joint region–class label table $\{p_{r,y}\}$.
The joint table is a reusable summary object: many operational KPIs can be written as linear projections (or monotone transforms) of $\{p_{r,y}\}$. Together with the induced Binomial/Beta–Binomial sampling model for events of interest, it provides finite-window predictive envelopes for future realized operational rates and a way to compare calibration settings without re-auditing. We also provide a conservative leave-one-out proxy for settings where there is no budget for an independent audit set (Appendix D).
@@
Together, these results clarify why some operational profiles are infeasible and help characterize the trade-offs induced by conformal calibration choices.
@@ 1.4 Positioning and related work
Conformal prediction constructs set-valued predictors with finite-sample coverage guarantees under exchangeability (Vovk et al., 2005; Shafer and Vovk, 2008; Angelopoulos and Bates, 2023). Conformal risk control (CRC) generalizes this perspective by choosing thresholds to satisfy user-specified scalar risk constraints under related assumptions (Angelopoulos et al., 2024; Bates et al., 2021), and extensions address covariate shift via reweighting or adaptation (Tibshirani et al., 2019; Fannjiang et al., 2022). Related work also studies stronger conditional notions of validity beyond marginal coverage, including training-conditional and other conditional guarantees (Bian and Barber, 2023; Gibbs et al., 2025). A growing literature connects conformal methods to downstream decision-making: conformal decision theory calibrates decision rules (often online) to control realized risk (Lekeufack et al., 2023), while decision-theoretic foundations motivate conformal sets as uncertainty objects for risk-averse agents (Kiyani et al., 2025).
Recent Human-Computer Interaction (HCI)-oriented work emphasizes a complementary concern: conformal sets can be a “murky” interface for human decision-making, because validity and common efficiency summaries do not identify the action-relevant consequences users care about (Hullman et al., 2025). Fig. 1 of our paper makes one such concern concrete: each conformal set corresponds to a region in the calibration-induced partition and is associated with specific region-conditional class frequencies, which are not apparent from set membership alone.
@@
Closest in spirit to our operational menu viewpoint is the inverse CRC of Zhou and Zhu (2025), who trace certified miscoverage–regret trade-offs across robustness levels for predict-then-optimize pipelines, auditing coverage and regret as the downstream objects of interest. A different line of work targets efficiency by selecting, from a family of candidate models, the conformal predictor that minimizes a set-size functional (e.g., interval length/volume, or prediction-set cardinality in classification) subject to validity constraints (Yang and Kuchibhotla, 2021). This is well-motivated when downstream utility is monotone in set size, but our objective differs: in many deployments, smaller cardinality does not necessarily imply higher utility, nor lower uncertainty in induced region-conditional label frequencies. In particular, aggressive singleton behavior can reduce average set size while increasing decisive error exposure and degrading the operational profile that stakeholders ultimately care about. When organizational priorities are still being determined, set size can therefore be an uninformative proxy for decision quality.
Our focus is complementary to these lines of work. Rather than guaranteeing coverage or a single scalar risk functional, we certify finite-window operational rate vectors for a fixed deployed rule and support multi-objective exploration without committing to a scalar objective. The audited primitive is the intermediate geometry induced by calibration—namely the region partition and its joint region–class label table—from which many deployment policies and KPIs follow by projection. This calibrate-and-audit scheme also clarifies the internal statistical structure of the prediction sets by making explicit how conformal calibration induces the partition from which downstream quantities are derived.
@@ 1.5 Roadmap
Section 2 formalizes the calibration-conditional viewpoint for a fixed deployed object and defines the region–label table as the auditable primitive. Section 3 presents SSBC and the Calibrate-and-Audit framework, yielding certified finite-window envelopes for operational rates and Pareto-based trade-off navigation. Section 4 characterizes regime boundaries and feasibility constraints induced by the conformal partition. Section 5 validates the framework empirically on Tox21 toxicity prediction and aqueous solubility prediction, and Section 6 discusses limitations and extensions.
@@ 2 Setting and notation: calibration-conditional viewpoint
We study a single deployed classifier produced by (i) training a scoring model on a training set and then treating it as fixed, followed by (ii) a one-time split conformal calibration on an exchangeable calibration sample. Accordingly, all randomness in our analysis is over the one-time calibration draw and future deployment data; we do not analyze variability due to retraining or repeated recalibration. Later, to make operational rates beyond coverage auditable with certified finite-sample guarantees, we introduce a separate exchangeable audit set held out alongside the calibration sample.
@@ 2.1 Data splits and exchangeability assumptions
Let $\mathcal{D}_{\mathrm{train}}$ denote a training dataset used to fit a base probabilistic classifier (Goodfellow et al., 2016). After training, the resulting score function is treated as fixed.
@@
The calibration sample $\mathcal{D}_{\mathrm{cal}}$ is used once to set the deployed rule’s thresholds; the audit sample $\mathcal{D}_{\mathrm{audit}}$ is held out and reserved for measuring operational quantities after thresholds are fixed, and is not used to choose thresholds. This use of a double calibration set is not without precedent; similar split designs appear in Yang and Kuchibhotla (2021).
@@ 2.2 Scores and calibration thresholds
Let $\mathcal{Y}=\{1,\ldots,K\}$, and let $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ denote a nonconformity (score) function (Shafer and Vovk, 2008; Lei et al., 2018). A common choice for probabilistic-like classifiers is $s(x,y)=1-P(y\mid x)$. While this specific form is not required, it is used in the examples throughout the paper.
@@
Thus, for a given $(\mathcal{D}_{\mathrm{cal}},\alpha^{\star})$, calibration produces a vector of class-conditional thresholds $\tau=(\tau_{1},\ldots,\tau_{K})$ that is held fixed during deployment.
@@ 2.3 Regions, observability, and policies
and class-conditional thresholds
@@
Thus calibration fixes a region map $R_{\tau}:\mathcal{X}\to\{0,1\}^{K}$, a finite partition of score space used throughout deployment. The map $R_{\tau}$ is determined entirely by $\tau$. However, the induced distribution of class labels $Y$ given the region $R_{\tau}(X)$, i.e. which regions carry mass and how much, depends on the deployment data-generating process. This discrete region partitioning is our decision and exploration interface.
Figure 2: Region definitions and region mass under different data regimes. (A) Thresholds $\tau_{0},\tau_{1}$ induce a finite region partition $R_{\tau}(x)$ independent of data and policy. (B–C) Under probability-normalized scores, data support is restricted to the diagonal, determining which regions carry nonzero mass for different threshold configurations. (D) With unconstrained scores, mass is distributed across a score-manifold, not solely no the diagonal. No deployment policy is applied in this figure.
@@
The standard split conformal predictor corresponds to the set inclusion policy:
@@ 2.4 Auditing primitives: region–label tables and policy projections
Conditional on $\mathcal{D}_{\mathrm{cal}}$, the deployed rule is fixed, and the primitive observable for auditing is the pair $(R_{\tau(\theta)}(X),Y)$ on an exchangeable labeled split. Here $\theta\in\Theta$ denotes a calibration setting that determines thresholds $\tau(\theta)$ and hence the region map $R_{\tau(\theta)}$. Two representations are useful: the full region–class label table, and operational indicators obtained as its projections.
(A) Region–class label joint table.
@@
The region map $R_{\tau(\theta)}$ is fixed from $\mathcal{D}_{\mathrm{cal}}$, while the counts $K^{\mathrm{audit}}_{r,y}(\theta)$ are computed on $\mathcal{D}_{\mathrm{audit}}$ only.
Because $R_{\tau(\theta)}$ is a finite partition of score space, the region–class label events $(R_{\tau(\theta)}(X)=r,\ Y=y)$ form a finite partition of the sample space. Consequently, for every $\theta$,
@@
Moreover, the class label marginal probabilities are fixed by the deployment data-generating process:
$\sum_{r\in\mathcal{R}}p_{r,y}(\theta)=\mathbb{P}(Y=y\mid\mathcal{D}_{\mathrm{cal}}),\qquad\sum_{r\in\mathcal{R}}\widehat{p}^{\,\mathrm{audit}}_{r,y}(\theta)=\widehat{\mathbb{P}}^{\,\mathrm{audit}}(Y=y).$
Varying $\theta$ reallocates mass across regions within each label. These conservation laws are the source of geometric coupling and explain why sweeping $\theta$ produces a constrained set of rate vectors rather than filling an ambient volume.
@@
Thus the region–label table is a reusable sufficient audit summary: audit it once, then compute policy-level KPIs by projection. The same structure extends to monotone transformations of linear KPIs, since the region–class label table remains the sufficient summary.
@@
$P(\theta):\begin{array}[]{c|ccc}&y=0&y=1&\hat{C}_{\pi_{\mathrm{SI}}}(X)\\ \hline\cr r_{10}=(1,0)&\mathbf{p_{10,0}(\theta)}&p_{10,1}(\theta)&\{0\}\\ r_{11}=(1,1)&\mathbf{p_{11,0}(\theta)}&\mathbf{p_{11,1}(\theta)}&\{0,1\}\\ r_{01}=(0,1)&p_{01,0}(\theta)&\mathbf{p_{01,1}(\theta)}&\{1\}\\ r_{00}=(0,0)&p_{00,0}(\theta)&p_{00,1}(\theta)&\varnothing\end{array}\qquad\text{with}\qquad\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}p_{r,y}(\theta)=1.$
@@
The same “sum selected cells” structure applies to abstention/deferral, decisiveness, and decisive error exposure under any fixed policy. In this sense, these operational quantities are all derived from a single auditable summary, from which many KPIs can be computed by projection while keeping the conservation constraints behind the trade-offs explicit.